OpenStudy (anonymous):

factor the expression 3 2x + 54 = 0 (the x is cubed)

OpenStudy (anonymous):

First we guess for one factor, I guessed and found that x-3 is a factor cause when x=3, the equation is equal to 0.

OpenStudy (anonymous):

x = -3 is the correct answer.

OpenStudy (anonymous):

I tried to get tutored and could not get help at school. I am not sure how yall got the answer.

OpenStudy (anonymous):

it isnt the final answer though u still have to use remainder factor theorem to get the other two factors

OpenStudy (anonymous):

\[2.x ^{3}+54 = 0\] \[2.x ^{3}= -54\] \[x ^{3} = -27\] \[x ^{3}= -3*-3*-3\] \[x^{3} = -3^{3}\] x = -3

OpenStudy (anonymous):

while sandeep gets the other factors i can help u to under stand this first part

OpenStudy (anonymous):

Hey Hawkson, just recheck, for x = 3 the equation doesnt satisfies.

OpenStudy (anonymous):

I appreciate the help. I am in pre algebra in college. our professor blew through this in 10 mins and I am so lost.

OpenStudy (anonymous):

Is that clear to you now!? Janedoepcola?

OpenStudy (anonymous):

but when i expand it i dont get back the original 2x^3 + 54 = 0 I get x^3 + 9x^2 + 27x + 27

OpenStudy (anonymous):

u understand it?

OpenStudy (anonymous):

Yeah, Because what you have done is \[(x+3)^{3}\] but actually it is in the form \[2*(x^{3}-(-3)^{3})=0\]

OpenStudy (anonymous):

When you factor out a 2, you end up with 2(x^3 +27)=0 You have to recognize x^3+27 as a sum of two cubes. The sum of two cubes factors as follows: \[x ^{3}+a ^{3}=(x+a)(x ^{2}-ax+a ^{2})\]